OpenStudy (hhlaine):

Which of these follows from Newton's laws about the energy of a moving object? A. the mechanical energy remains the same if there are only conservative forces acting. B. the total mechanical energy of an object is always constant. C. the kinetic energy of an object remains the same if friction is absent. D. the total all forms of energy is always constant.

OpenStudy (shootingstar28):

I think the answer could be A, the reason why is because we can write the subsequent chain of equalities, for the free falling particle whose mass is m: PE=mgh=mg1/2m(gt)^2=1/2mv^2=KE

OpenStudy (hhlaine):

@ShootingStar28 oooooh okay that makes sense. Thank you!! I have another one I'm stuck on, would you mind trying to help?

rebeccaxhawaii (rebeccaxhawaii):

\(\Huge\boxed{\boxed{\boxed{\huge\mathbb{WELCOME~TO~OPENSTUDY} }}}\) Hey i have made this to hopefully answer any questions you may have. But if it doesn't feel free to pm me c: http://openstudy.com/users/rebeccaxhawaii#/updates/56f5205ce4b07a8c82287f75 Good Luck with you studies! \(\Huge\heartsuit\)